OpenStudy (anonymous):

3.2 x 10-6/8 x 10-4 the x is a multiplication sign. the -6 and -4 are exponents. the / is a division sign.

OpenStudy (amistre64):

3.2 (10^-6) ---------- like this? 8 (10^-4)

OpenStudy (amistre64):

or is 10-6 mean 10^6?

OpenStudy (anonymous):

3.2 times 10 to the power of 8 divided by 7 times 10 to the power of 4

OpenStudy (amistre64):

.... just for my sanity, tell me again what the problem states :) cause youve got 2 different things here now

OpenStudy (anonymous):

i got the answer. heres another question: \[3.5\times10^{8}\div7\times10^{4}\]

OpenStudy (amistre64):

ok... I see that 35/70 = 1/2 or .5 lets go with .5 .5 x 10^(8-4) = .5 (10^4) to keep things in scietific notation, we should move the decimal till we get an integer like this: 5 (10^5)

OpenStudy (amistre64):

might be 5 (10^3) i cant remember which way that goes :)

OpenStudy (amistre64):

its 5 x 10^3

OpenStudy (anonymous):

heres another question: \[(3xy ^{3})^{-2}\div (x ^{4}y)^{-2}\]